OpenStudy (anonymous):

In radical form, 81^1/3 is written as ___ or ___.

OpenStudy (anonymous):

In radical form, \[81^{1/3} \] is written as ___ or ___.

OpenStudy (anonymous):

One answer is 729 right? So, tell me the other pls?

OpenStudy (anonymous):

Why would 729 be one answer?

OpenStudy (shamim):

4.33

OpenStudy (anonymous):

Don't just give the answer @shamim

OpenStudy (shamim):

\[81^{1/3}=4.33\]

OpenStudy (anonymous):

Help her learn.

OpenStudy (anonymous):

@shamim there's no other fraction form?

OpenStudy (anonymous):

sorry. i meant \[81^{3/2}\]

OpenStudy (shamim):

ok

OpenStudy (shamim):

\[81^{3/2}=729\]

OpenStudy (anonymous):

Yup:) 729. what's the other?

OpenStudy (shamim):

it has only 1 answer

OpenStudy (shamim):

as i know

OpenStudy (anonymous):

The question asks: "In radical form, (81)^(2/3) is written as ___ or ___."

OpenStudy (anonymous):

or can be 729/1?

OpenStudy (anonymous):

Do you know what "radical form" is?

OpenStudy (anonymous):

"Radical" form is a form that invovles the radical symbol, right?

OpenStudy (anonymous):

hmm yeah gaaad its hard ><

OpenStudy (anonymous):

\[\sqrt[n]{m}\]

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sqrt[3/2]{81}\] ?